A metal sphere of radius 'r' and specific heat  's' is rotated about an axis paying through is centre at a speed of 'n'  rotations. If it suddenly stops and 50% of its energy is used in increasing its temperature. The rise in temperature of sphere is

Grade:12

1 Answers

Arun
5 years ago
A.T.Q 
It's rotational K.E. = n*m*(r^2)*pi / 5
So 50% energy = n*m*(r^2)*pi / 10
So that much energy is absorbed by sphere to increase it's temperature:
So
m*S*(del. T) = n*m*(r^2)*pi / 10
So del. T = n * r^2 *pi / (10*S)